OpenStudy (anonymous):

solve and show steps 3root(x-2) -8=8

OpenStudy (anonymous):

is -8 under the root ??

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Add 8 to both sides. 3 root (x-2) = 16 divide by 3 root (x-2) = 16 / 3 square both sides x - 2 = 256 / 9 add 2 x = 274 / 9

OpenStudy (lalaly):

\[3\sqrt{x-2}=16\] \[\sqrt{x-2}= \frac{16}{3}\] \[x-2 = (\frac{16}{3})\] now add 2 to both sides

OpenStudy (lalaly):

sorry (16/3)^2